# Hallucination Detection in Large Language Models with Metamorphic Relations

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3715735conference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYjournal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE020publicationmonth: 7ccs: Computing methodologies Natural language processingccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Large Language Models (LLMs) like GPT-4 (35) have revolutionized language processing by generating fluent and organized responses for various applications, such as drafting reports and summarization systems (34; 24; 54). However, LLMs are prone to generating hallucinations—coherent but factually incorrect or irrelevant outputs, such as non-factual responses in question-answering context (17; 47). This tendency poses significant challenges to the reliability of LLMs, undermining the effectiveness of LLMs in applications requiring high factual accuracy. Fact-conflicting hallucinations, where LLMs produce content that contradicts established facts, are particularly concerning as they can mislead users lacking expertise on the topic, leading to significant confusion and eroding the trust essential for various LLM applications. Figure 1 illustrates a scenario where a legal question is posed to ChatGPT, which generates a response containing hallucinations. Without proper detection, such inaccuracies could have serious consequences for individuals who lack expertise in the legal field.
@@
Many approaches have been introduced to detect or mitigate hallucination of LLMs that compare responses to factual information, often relying on databases or search engines (15; 7; 16; 6; 14; 18). However, relying on external resources for hallucination detection often limits the scope to specific domains where a comprehensive database does not exist. Moreover, hallucinations are observed across a wide range of tasks that extend beyond simple fact verification (23). Other works use token-level information such as token confidence and entropy (43; 49), which is often inaccessible in closed-source models. To tackle these issues, Manakul et al. proposed SelfcheckGPT (31), a self-contained approach to addressing hallucination. However, for the sample generation process in SelfCheckGPT, the LLM tends to produce samples that are identical to the original response, which significantly impacts its performance in hallucination detection.
@@
    We improve the TruthfulQA (29) benchmark by updating 238 questions with new correct answers. We share the improved benchmark, which we name TruthfulQA-Enhanced. This new benchmark can support more accurate hallucination detection research.
@@ 2.1. Motivating Example
In contrast, MetaQA applies Metamorphic Relations (MR) (9; 52) to introduce controlled mutations by generating a set of mutations from the base response. Formally, let $B$ be the base response. MetaQA defines a function $f(B)$ using the same LLM that produces two sets of mutations: a set of synonymous mutations $S=f_{s}(B)=\{s_{0},s_{1},s_{2},\dots,s_{N}\}$ and a set of antonymous mutations $A=f_{a}(B)=\{a_{0},a_{1},a_{2},\dots,a_{M}\}$. For each mutation, the system verifies whether the transformation remains factually consistent. This verification process assigns a hallucination score based on the factual alignment of both mutation sets $S$ and $A$. The advantage of MR over repeated prompts is that MR introduces changes to the input that make LLMs reveal inconsistencies more effectively, particularly when hallucinations are present (19). While repeated prompts often produce consistent outputs—even if the base response contains hallucinations—MR’s varied mutations expose these inconsistencies by prompting the LLM to produce divergent responses. Additionally, independently verifying each mutation reduces bias from previous outputs, resulting in a more accurate and reliable detection of hallucinations (12). Revisiting the example shown in Figure 2, we observe that while SelfCheckGPT computed a very low hallucination score, our approach—which uses various MRs and validates each mutation individually—resulted in a high hallucination score of $1.0$ as none of the mutations were validated as a fact. This indicates a greater likelihood of detecting hallucinations in this case.
@@ 2.3. LLM Hallucination Types
Hallucination of LLMs has been extensively studied in the realm of Natural Language Processing (NLP). Hallucination typically refers to a phenomenon where the generated content appears nonsensical or unfaithful to the provided source content (20). Generally, hallucination in natural language generation tasks can be categorized into three primary types (50; 55), as detailed below:
  • Input-Conflicting Hallucination: This type occurs when LLM generates outputs that are inconsistent with the user’s input. Typically, such inconsistencies can present themselves in two primary ways: the LLM’s response might conflict with the user’s task instructions, suggesting a misinterpretation of the intent, or the generated output might contradict the task input, similar to common issues in machine translation, or summarization (27).
@@
  • Fact-Conflicting Hallucination occur when generated content contradicts established world knowledge (8; 10; 32). Several factors throughout the life-cycle of an LLM, including the training dataset, pre-training procedures, and the inference process, can contribute to such type of hallucination. For instance, as shown in Figure 1, an LLM might present inaccurate information in response to a user query, potentially misleading users who lack expertise on the topic.
@@ 2.4. Metamorphic Relation (MR)
A metamorphic relationship $R$ is a necessary property between the set of inputs $X=<x_{1},x_{2},...,x_{n}>$ and their corresponding outputs $Y=<f(x_{1}),f(x_{2}),...f(x_{n})>$ from a given function $f$ and is denoted as $R\subseteq X^{n}\times Y^{n}$ or simply as $R(x_{1},x_{2},...,x_{n},f(x_{1}),f(x_{2}),...f(x_{n}))$ (9; 52). Metamorphic relations can be applied independently or in conjunction with other static and dynamic software analysis techniques, such as formal proofs and debugging. Essentially, an MR is a property or constraint that the output of a system should satisfy when specific transformations or modifications are applied to its input. If the system’s output changes inappropriately in response to these transformations, it may indicate that the system is producing incorrect or unreliable results.
@@ 3. MetaQA Methodology
We guide LLM to produce concise and fact-based answers in response to an input question. In typical interactions, LLMs often generate detailed and lengthy explanations. However, within the MetaQA framework, excessive verbosity can hinder downstream modules involved in tasks such as mutation generation and verification. To address this, we employ system prompts that instruct the LLM to generate brief, contextually grounded answers. This ensures that responses remain precise. Recent studies have shown that LLMs are effective in evaluating the consistency of information between lengthy documents and concise summaries, even in zero-shot settings (30). Therefore, in addition to the standard question-answering process, we instruct the model to summarize its responses into short, accurate sentences.
@@
For a given response (we call it base response) to a question from Step 1, we create multiple distinct high-quality mutations to the response. Each mutation is produced as a follow-up question to the base response. MetaQA employs a prompt-based approach that exclusively uses the LLM to generate mutations111 We do not adopt traditional synonym- and antonym-based mutation generation methods (40; 53) because LLMs have consistently demonstrated superior performance over traditional NLP techniques across almost all NLP tasks (37; 13).. Thus, for potentially overly brief responses from the LLM, this method will generate complete and semantically accurate mutations based on the context of the question. A mutation is denoted as being generated in the following manner:
@@
We deploy two general types of reasoning rules prevalently adopted in several literature (3; 28; 38; 41; 56), the types of metamorphic relations are detailed as follows:
@@ 4.1. Baseline
We use SelfCheckGPT (31), the state-of-the-art (SOTA) hallucination detection approach that does not need external resources. As outlined in Section 2.1, by repeatedly querying an LLM to generate reference samples and by measuring their consistency with the original response, SelfCheckGPT calculates a hallucination score for user reference. As such, both MetaQA and SelfCheckGPT use a threshold based on Equation 5) to determine whether a response is hallucinated or not. We used the publicly available version of SelfCheckGPT (31), which by default calls the ChatGPT API without specifying an explicit temperature value. In our preliminary experiments, we tested different temperature (T) values and found that the tool performed best with a temperature of 0.5 (specific results with different temperatures are shown on our homepage [^2] https://github.com/zbybr/LLMhalu/tree/ForMetaQAPaper). Therefore, in our evaluation, by default, we use threshold 0.5 to compare MetaQA and SelfCheckGPT. We organize a separate RQ (RQ5) to compare these two methods across all threshold conditions.
@@ 4.2. Datasets
We adopt three benchmark datasets: TruthfulQA (29), HotpotQA (48), and FreshQA (44). Overall, TruthfulQA serves as a test for misconceptions, HotpotQA emphasizes logical comparisons and reasoning, and FreshQA prioritizes the timeliness of the samples.
@@ 4.2.1. TruthfulQA
We use TruthfulQA (29) as our primary experimental dataset. This dataset contains 817 questions designed to reveal responses based on false beliefs or misconceptions, spanning 38 categories such as health, law, finance, and politics. During our initial experiments, we discovered that many potentially correct LLM responses were not included in the ‘correct answers’ provided by TruthfulQA. Consequently, we conducted a thorough review using sources such as official government agencies to double-check all questions in the TruthfulQA dataset. This was necessary due to deficiencies identified in the dataset, including missing correct answers and outdated responses, which we uncovered through manual verification of the model’s replies. As a result, we updated 238 entries in the TruthfulQA dataset. These updates were crucial, as they provided more accurate answers to many questions over time. We refer to this revised dataset as ‘TruthfulQA-Enhanced’, a sample is shown in Table 2.
@@ 4.2.2. HotpotQA
HotpotQA (48) is also part of our experimental dataset. HotpotQA is a question-answering dataset featuring natural, multi-hop questions with strong supervision for supporting facts, facilitating more interpretable question-answering systems. The dataset covers a wide range of real-world domains and comprises a total of 113K questions, making it an excellent supplement to TruthfulQA. Many questions in HotpotQA involve comparative reasoning across two or more items, providing a robust test for the reasoning capabilities of LLMs and allowing us to observe hallucinations in questions requiring logical reasoning. However, the large number of question instances in HotpotQA and the often overly simplistic reference answers can lead to increased token consumption by MetaQA for summary expansion and answer comparison. To address this issue, we randomly selected 610 questions from HotpotQA.
@@ 4.2.3. FreshQA
We include FreshQA (44), which features questions requiring up-to-date world knowledge and those based on false premises. To ensure a fair comparison, we selected 155 questions from FreshQA dated before 2023, as LLMs may generate hallucinations when addressing newer questions due to potential gaps in training data. Overall, our experimental dataset encompasses 1582 questions across multiple datasets.
@@ 4.3. Studied LLMs
To ensure a reliable evaluation in our experiments, we assess several LLMs that have been evaluated in LiveBench (45), a benchmark used to assess LLMs across various aspects, including math, reasoning, language, instruction following, and more. Our evaluation includes very large models from the GPT family as well as open-source models from the Llama and Mistral families. We divide the selected models into two types:
@@ 5.3.1. Stability over multiple runs
Previous results demonstrate that MetaQA is an effective approach for detecting hallucinations in the question-answering process with LLMs. An additional consideration is the stability of MetaQA as a hallucination detection method. Given the randomness of hallucination generation in LLMs (36) and the reliance of MetaQA’s workflow on the LLM itself, it is crucial to assess whether MetaQA is affected by these uncertainties. To address this concern, we repeated MetaQA runs 3 times on the same dataset using both the open-source model Llama3 and the closed-source model GPT-3.5.
@@ 5.3.2. Impact of temperature
Temperature plays a crucial role in the functioning of LLMs by influencing the randomness and creativity of the generated text (51). In LLMs, temperature serves as a parameter that controls output diversity by manipulating the softmax function, which determines the probabilities of the next word in a sequence. When the temperature is low (e.g., close to zero), the generated text tends to be more focused and deterministic, leading to more confident predictions and less variation in the output. This characteristic can be beneficial when a conservative and predictable response is preferred. For MetaQA, accurate mutation generation and verification results are essential, making it imperative to investigate the impact of temperature on its performance. In our experiments, we assessed the stability of MetaQA across a set of temperature values, denoted as $T=\{0.1,0.3,0.5,0.7\}$, which were selected as the experimental parameters. We randomly sampled 10% of the total dataset for the temperature experiments conducted on GPT-3.5. As illustrated in Figure 8, this finding aligns with our anticipated results, demonstrating that MetaQA’s performance decreases with increasing temperatures and performs better at lower temperatures.
@@ 6.2. Importance of Hallucination Detection
Hallucination detection is similar to software testing and machine learning testing (53), where bug detection, though expected, is critical for improving system reliability and building user trust. Similarly, detecting hallucinations in LLMs helps establish trust boundaries, refine the model design, and ensure accuracy in applications in high-stakes fields(17; 47). While hallucinations, like bugs, cannot be fully eliminated, their detection and management allow us to mitigate risks and enhance the reliability of LLMs. Given their impact on trust and safety, hallucination detection deserves dedicated effort and innovation. We call on the community to contribute to advancing robust hallucination detection methods, improving mitigation strategies, to ensuring the responsible deployment of LLM systems.

### v2

@@ Abstract.
copyright: acmlicenseddoi: 10.1145/3715735journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE020publicationmonth: 7conference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYccs: Computing methodologies Natural language processingccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Large Language Models (LLMs) like GPT-4 (OpenAI et al., 2023) have revolutionized language processing by generating fluent and organized responses for various applications, such as drafting reports and summarization systems (Moramarco et al., 2021; Lai et al., 2022; Zhang et al., 2023a). However, LLMs are prone to generating hallucinations—coherent but factually incorrect or irrelevant outputs, such as non-factual responses in question-answering context (Huang et al., 2023; Xu et al., 2024). This tendency poses significant challenges to the reliability of LLMs, undermining the effectiveness of LLMs in applications requiring high factual accuracy. Fact-conflicting hallucinations, where LLMs produce content that contradicts established facts, are particularly concerning as they can mislead users lacking expertise on the topic, leading to significant confusion and eroding the trust essential for various LLM applications. Figure 1 illustrates a scenario where a legal question is posed to ChatGPT, which generates a response containing hallucinations. Without proper detection, such inaccuracies could have serious consequences for individuals who lack expertise in the legal field.
@@
Many approaches have been introduced to detect or mitigate hallucination of LLMs that compare responses to factual information, often relying on databases or search engines (Guo et al., 2022; Augenstein et al., 2019; Hanselowski et al., 2019; Atanasova et al., 2020; Galitsky, 2023; Huo et al., 2023). However, relying on external resources for hallucination detection often limits the scope to specific domains where a comprehensive database does not exist. Moreover, hallucinations are observed across a wide range of tasks that extend beyond simple fact verification (Kryściński et al., 2019). Other works use token-level information such as token confidence and entropy (Varshney et al., 2023; Yao et al., 2023), which is often inaccessible in closed-source models. To tackle these issues, Manakul et al. proposed SelfcheckGPT (Manakul et al., 2023), a self-contained approach to addressing hallucination. However, for the sample generation process in SelfCheckGPT, the LLM tends to produce samples that are identical to the original response, which significantly impacts its performance in hallucination detection.
@@
    We improve the TruthfulQA (Lin et al., 2021) benchmark by updating 238 questions with new correct answers. We share the improved benchmark, which we name TruthfulQA-Enhanced. This new benchmark can support more accurate hallucination detection research.
@@ 2.1. Motivating Example
In contrast, MetaQA applies Metamorphic Relations (MR) (Chen et al., 2018; Zhang et al., 2014) to introduce controlled mutations by generating a set of mutations from the base response. Formally, let $B$ be the base response. MetaQA defines a function $f(B)$ using the same LLM that produces two sets of mutations: a set of synonymous mutations $S=f_{s}(B)=\{s_{0},s_{1},s_{2},\dots,s_{N}\}$ and a set of antonymous mutations $A=f_{a}(B)=\{a_{0},a_{1},a_{2},\dots,a_{M}\}$. For each mutation, the system verifies whether the transformation remains factually consistent. This verification process assigns a hallucination score based on the factual alignment of both mutation sets $S$ and $A$. The advantage of MR over repeated prompts is that MR introduces changes to the input that make LLMs reveal inconsistencies more effectively, particularly when hallucinations are present (Hyun et al., 2024). While repeated prompts often produce consistent outputs—even if the base response contains hallucinations—MR’s varied mutations expose these inconsistencies by prompting the LLM to produce divergent responses. Additionally, independently verifying each mutation reduces bias from previous outputs, resulting in a more accurate and reliable detection of hallucinations (Dhuliawala et al., 2024). Revisiting the example shown in Figure 2, we observe that while SelfCheckGPT computed a very low hallucination score, our approach—which uses various MRs and validates each mutation individually—resulted in a high hallucination score of $1.0$ as none of the mutations were validated as a fact. This indicates a greater likelihood of detecting hallucinations in this case.
@@ 2.3. LLM Hallucination Types
Hallucination of LLMs has been extensively studied in the realm of Natural Language Processing (NLP). Hallucination typically refers to a phenomenon where the generated content appears nonsensical or unfaithful to the provided source content (Ji et al., 2023). Generally, hallucination in natural language generation tasks can be categorized into three primary types (Yao et al., 2024; Zhang et al., 2023b), as detailed below:
  • Input-Conflicting Hallucination: This type occurs when LLM generates outputs that are inconsistent with the user’s input. Typically, such inconsistencies can present themselves in two primary ways: the LLM’s response might conflict with the user’s task instructions, suggesting a misinterpretation of the intent, or the generated output might contradict the task input, similar to common issues in machine translation, or summarization (Li et al., 2022).
@@
  • Fact-Conflicting Hallucination occur when generated content contradicts established world knowledge (Chen et al., 2023; Chern et al., 2023; Min et al., 2023). Several factors throughout the life-cycle of an LLM, including the training dataset, pre-training procedures, and the inference process, can contribute to such type of hallucination. For instance, as shown in Figure 1, an LLM might present inaccurate information in response to a user query, potentially misleading users who lack expertise on the topic.
@@ 2.4. Metamorphic Relation (MR)
A metamorphic relationship $R$ is a necessary property between the set of inputs $X=<x_{1},x_{2},...,x_{n}>$ and their corresponding outputs $Y=<f(x_{1}),f(x_{2}),...f(x_{n})>$ from a given function $f$ and is denoted as $R\subseteq X^{n}\times Y^{n}$ or simply as $R(x_{1},x_{2},...,x_{n},f(x_{1}),f(x_{2}),...f(x_{n}))$ (Chen et al., 2018; Zhang et al., 2014). Metamorphic relations can be applied independently or in conjunction with other static and dynamic software analysis techniques, such as formal proofs and debugging. Essentially, an MR is a property or constraint that the output of a system should satisfy when specific transformations or modifications are applied to its input. If the system’s output changes inappropriately in response to these transformations, it may indicate that the system is producing incorrect or unreliable results.
@@ 3. MetaQA Methodology
We guide LLM to produce concise and fact-based answers in response to an input question. In typical interactions, LLMs often generate detailed and lengthy explanations. However, within the MetaQA framework, excessive verbosity can hinder downstream modules involved in tasks such as mutation generation and verification. To address this, we employ system prompts that instruct the LLM to generate brief, contextually grounded answers. This ensures that responses remain precise. Recent studies have shown that LLMs are effective in evaluating the consistency of information between lengthy documents and concise summaries, even in zero-shot settings (Luo et al., 2023). Therefore, in addition to the standard question-answering process, we instruct the model to summarize its responses into short, accurate sentences.
@@
For a given response (we call it base response) to a question from Step 1, we create multiple distinct high-quality mutations to the response. Each mutation is produced as a follow-up question to the base response. MetaQA employs a prompt-based approach that exclusively uses the LLM to generate mutations111 We do not adopt traditional synonym- and antonym-based mutation generation methods (Sun et al., 2020; Zhang et al., 2022) because LLMs have consistently demonstrated superior performance over traditional NLP techniques across almost all NLP tasks (Qin et al., 2024; Foster et al., 2025).. Thus, for potentially overly brief responses from the LLM, this method will generate complete and semantically accurate mutations based on the context of the question. A mutation is denoted as being generated in the following manner:
@@
We deploy two general types of reasoning rules prevalently adopted in several literature (Abboud et al., 2020; Liang et al., 2024; Ren and Leskovec, 2020; Tian et al., 2022; Zhou et al., 2019), the types of metamorphic relations are detailed as follows:
@@ 4.1. Baseline
We use SelfCheckGPT (Manakul et al., 2023), the state-of-the-art (SOTA) hallucination detection approach that does not need external resources. As outlined in Section 2.1, by repeatedly querying an LLM to generate reference samples and by measuring their consistency with the original response, SelfCheckGPT calculates a hallucination score for user reference. As such, both MetaQA and SelfCheckGPT use a threshold based on Equation 5) to determine whether a response is hallucinated or not. We used the publicly available version of SelfCheckGPT (Manakul et al., 2023), which by default calls the ChatGPT API without specifying an explicit temperature value. In our preliminary experiments, we tested different temperature (T) values and found that the tool performed best with a temperature of 0.5 (specific results with different temperatures are shown on our homepage [^2] https://github.com/zbybr/LLMhalu/tree/ForMetaQAPaper). Therefore, in our evaluation, by default, we use threshold 0.5 to compare MetaQA and SelfCheckGPT. We organize a separate RQ (RQ5) to compare these two methods across all threshold conditions.
@@ 4.2. Datasets
We adopt three benchmark datasets: TruthfulQA (Lin et al., 2021), HotpotQA (Yang et al., 2018), and FreshQA (Vu et al., 2023). Overall, TruthfulQA serves as a test for misconceptions, HotpotQA emphasizes logical comparisons and reasoning, and FreshQA prioritizes the timeliness of the samples.
@@ 4.2.1. TruthfulQA
We use TruthfulQA (Lin et al., 2021) as our primary experimental dataset. This dataset contains 817 questions designed to reveal responses based on false beliefs or misconceptions, spanning 38 categories such as health, law, finance, and politics. During our initial experiments, we discovered that many potentially correct LLM responses were not included in the ‘correct answers’ provided by TruthfulQA. Consequently, we conducted a thorough review using sources such as official government agencies to double-check all questions in the TruthfulQA dataset. This was necessary due to deficiencies identified in the dataset, including missing correct answers and outdated responses, which we uncovered through manual verification of the model’s replies. As a result, we updated 238 entries in the TruthfulQA dataset. These updates were crucial, as they provided more accurate answers to many questions over time. We refer to this revised dataset as ‘TruthfulQA-Enhanced’, a sample is shown in Table 2.
@@ 4.2.2. HotpotQA
HotpotQA (Yang et al., 2018) is also part of our experimental dataset. HotpotQA is a question-answering dataset featuring natural, multi-hop questions with strong supervision for supporting facts, facilitating more interpretable question-answering systems. The dataset covers a wide range of real-world domains and comprises a total of 113K questions, making it an excellent supplement to TruthfulQA. Many questions in HotpotQA involve comparative reasoning across two or more items, providing a robust test for the reasoning capabilities of LLMs and allowing us to observe hallucinations in questions requiring logical reasoning. However, the large number of question instances in HotpotQA and the often overly simplistic reference answers can lead to increased token consumption by MetaQA for summary expansion and answer comparison. To address this issue, we randomly selected 610 questions from HotpotQA.
@@ 4.2.3. FreshQA
We include FreshQA (Vu et al., 2023), which features questions requiring up-to-date world knowledge and those based on false premises. To ensure a fair comparison, we selected 155 questions from FreshQA dated before 2023, as LLMs may generate hallucinations when addressing newer questions due to potential gaps in training data. Overall, our experimental dataset encompasses 1582 questions across multiple datasets.
@@ 4.3. Studied LLMs
To ensure a reliable evaluation in our experiments, we assess several LLMs that have been evaluated in LiveBench (white2024livebenchchallengingcontaminationfreellm), a benchmark used to assess LLMs across various aspects, including math, reasoning, language, instruction following, and more. Our evaluation includes very large models from the GPT family as well as open-source models from the Llama and Mistral families. We divide the selected models into two types:
@@ 5.3.1. Stability Over Multiple Runs
Previous results demonstrate that MetaQA is an effective approach for detecting hallucinations in the question-answering process with LLMs. An additional consideration is the stability of MetaQA as a hallucination detection method. Given the randomness of hallucination generation in LLMs (Ouyang et al., 2025) and the reliance of MetaQA’s workflow on the LLM itself, it is crucial to assess whether MetaQA is affected by these uncertainties. To address this concern, we repeated MetaQA runs 3 times on the same dataset using both the open-source model Llama3 and the closed-source model GPT-3.5.
@@ 5.3.2. Impact Of Temperature
Temperature plays a crucial role in the functioning of LLMs by influencing the randomness and creativity of the generated text (Yu et al., [n. d.]). In LLMs, temperature serves as a parameter that controls output diversity by manipulating the softmax function, which determines the probabilities of the next word in a sequence. When the temperature is low (e.g., close to zero), the generated text tends to be more focused and deterministic, leading to more confident predictions and less variation in the output. This characteristic can be beneficial when a conservative and predictable response is preferred. For MetaQA, accurate mutation generation and verification results are essential, making it imperative to investigate the impact of temperature on its performance. In our experiments, we assessed the stability of MetaQA across a set of temperature values, denoted as $T=\{0.1,0.3,0.5,0.7\}$, which were selected as the experimental parameters. We randomly sampled 10% of the total dataset for the temperature experiments conducted on GPT-3.5. As illustrated in Figure 8, this finding aligns with our anticipated results, demonstrating that MetaQA’s performance decreases with increasing temperatures and performs better at lower temperatures.
@@ 6.2. Importance of Hallucination Detection
Hallucination detection is similar to software testing and machine learning testing (Zhang et al., 2022), where bug detection, though expected, is critical for improving system reliability and building user trust. Similarly, detecting hallucinations in LLMs helps establish trust boundaries, refine the model design, and ensure accuracy in applications in high-stakes fields(Huang et al., 2023; Xu et al., 2024). While hallucinations, like bugs, cannot be fully eliminated, their detection and management allow us to mitigate risks and enhance the reliability of LLMs. Given their impact on trust and safety, hallucination detection deserves dedicated effort and innovation. We call on the community to contribute to advancing robust hallucination detection methods, improving mitigation strategies, to ensuring the responsible deployment of LLM systems.